the proportion of students in private schools is around 11%. A random sample of 450 students from a wide geographic area indicated that 55 attended private schools. Estimate the true proportion of students attending private schools with 95% confidence. How does your estimate compare to 11%?
Answer by stanbon(75887)

the proportion of students in private schools is around 11%. A random sample of 450 students from a wide geographic area indicated that 55 attended private schools. Estimate the true proportion of students attending private schools with 95% confidence. ---
p-hat = 55/450 = 0.1222
---
E = 1.96*sqrt[0.11*0.89/450] = 0.0288
===================================
95% CI: 0.1222- 0.0288 < p < 0.1222+ 0.0288
95% CI: 0.934 < p < 0.1510
----------------------------------
How does your estimate compare to 11%?
The confidence inteval contains 11%
==============================================
Cheers,
Stan H.